Show that the sum of three consecutive integers is always equal to 3 times the middle integer

Answer by Edwin McCravy(20054)

smallest integer = n
middle integer = n+1
largest integer = n+2

Their sum = n+(n+1)+(n+2) = n+n+1+n+2 = 3n+3 = 3(n+1)

3(n+1) is 3 times n+1, which is the middle integer.

So we have proved it.

Edwin